# From Parameters to Behaviors: Unsupervised Compression of the Policy Space

### v1

@@ 5.2 Supervised Fine-Tuning
(Q3) How can we fine-tune against specific tasks, leveraging the low-dimensional space? Does this come with positives?
To achieve this, we compared the effect of supervised fine-tuning on the latent space with various baselines, including PPO (Schulman et al., 2017), DDPG (Lillicrap et al., 2015), TD3 (Fujimoto et al., 2018), and SAC (Haarnoja et al., 2018). We also tested PGPE (Sehnke et al., 2008) in the high-dimensional parameter space, referred to as Parameter PGPE, as a sanity check. All the algorithms were run in the best-performing configuration for the policy sizes. Interestingly, DRL baselines always struggle to optimize small policies and typically perform better with larger ones. On the contrary, Parameter PGPE benefits from having a reduced set of parameters to control; however, it generally suffers from high sample complexity. These evidences are reported in Appendix B (Fig. 6,7,8).
From these comparisons (reported in Fig. 4 and Fig. 5 for MC and RC, respectively), we were able to extract three main findings. First of all, the convergence rate and performance of Latent PGPE are positively correlated with the number of dimensions of the latent space, which confirms that larger latent spaces are indeed better shaped and with an easier optimization landscape, as hinted in Subsec. 5.1 as well. Secondly, Latent PGPE converges faster than all baselines in 7 out of 8 tasks, even though it does not always converge to the optimum. Finally, Latent PGPE is able to achieve comparable, if not better, performances to most of the baselines, even for complex tasks like speed in Fig. 4(c), 5(a). However, we observe that it fails of solve the height task in Fig. 4(d), due to the scarce representation of the high-performance policies in the unsupervised policy dataset.
Takeaways.  These results provide a positive answer to (Q3): Leveraging the learned low-dimensional representation of the behavioral manifold, the agent can not only achieve faster convergence, but also better performances than state-of-the-art DRL algorithms in challenging sparse tasks. Unfortunately, this comes with a bitter lesson: the fine-tuning performance is related to the quality of the learned representation and how the policy dataset is effective in learning a rich latent representation.
@@
Figure 5: Performance comparison in RC for different tasks. We report the average and 95% confidence interval over 10 runs. For clarity, the worst performing baselines are omitted. A full study is reported in Appendix B.
@@ 6 Conclusions
In this work, we proposed a novel, unsupervised framework to address the sample inefficiency of Deep Reinforcement Learning by shifting the focus from parameter space to behavior space. Our approach successfully learns a compact latent manifold of policies, organized by behavioral similarity, using a generative model with an unsupervised behavioral reconstruction loss. Empirically, we showed that this approach can compress policy parameterizations by several orders of magnitude while preserving their functional expressivity. This compressed representation also allows for more efficient fine-tuning for downstream tasks via gradient-based optimization in the low-dimensional latent space.
@@
All experiments were run on 50 cores of an Intel(R) Xeon(R) Gold 64118H CPU, 1 TB of RAM, and one NVIDIA H100 GPU. The total wall-clock time to re-run all experiments is approximately 50 hours. Our proposed approach is detailed in section with full implementation details and hyperparameters included in Appendix B. The source code is available in the supplementary material.
@@ 7 Acknowledgments
This work was supported by the Italian Ministry of University and Research (MUR) under the National Recovery and Resilience Plan (NRRP), and by the European Union (EU) under the NextGenerationEU project (CUP D43C24001670008)
Appendix A Related Work
@@
Simplification of the Policy Space.  A variety of works (among others, Gregor et al., 2017; Eysenbach et al., 2018; Achiam et al., 2018; Hansen et al., 2019) have proposed methods to simplify the policy space. Yet, those policies should be generally intended as mere initializations for supervised fine-tuning, which falls back to operating in the original policy space once the downstream task is revealed. To the best of our knowledge, the only other work defining a formal criterion to operate a compression of the policy space is Mutti et al. (2022). Yet, this paper seeks a way to reduce the cardinality of the policy space, rather than its dimensionality. Moreover, the constraints defining a valid compression are stricter than ours, resulting in an optimization problem that is known to be NP-hard. Finally, their work do not provide a way to perform supervised fine-tuning in a scalable way.
@@
Manifolds and Latent Representations of Policies.  The idea of employing generative models to learn a compressed representation of the policy space has received some recent attentions, and Rakicevic et al. (2021) hypothesized that there might be a low-dimensional manifold embedded in the policy parameter space, even if they did not characterize it formally. Chang et al. (2019) trains a Variational AE to reconstruct the weights of pre-trained expert policies in order to learn expert-agent embeddings and analyzing the latent structure of the solution space. A similar architecture has also been applied in the field of Quality Diversity, either to improve the sample efficiency of diversity-based search Rakicevic et al. (2021) or to distill a large policy archive into a compact generative model  (Hegde et al., 2023). Notably, all the methods above employ VAE architectures with a parameter-reconstruction loss, which allows only moderate compression ratios of up to $19:1$ (Hegde et al., 2023). In comparison, this paper introduces a fully unsupervised pipeline that focuses on compressing a behavioral loss, intending to provide a compact space for latent policy optimization as well while retaining way stronger compression abilities.
@@
Policy Optimization.  First-order methods have been extensively employed to address PO (Peters & Schaal, 2008; Lillicrap et al., 2015) as well as natural gradients (Kakade, 2001) and trust-region methods (Schulman et al., 2017). Yet in this work, we built over the long tradition of PGPE Algorithms (Sehnke et al., 2008; Rückstiess et al., 2010; Miyamae et al., 2010; Montenegro et al., 2024), as their hard scalability to large parameter spaces is notoriously a blocking factor. Finally, we notice that Rakicevic et al. (2021) indeed proposed a method to optimize the diversity of the policies by taking into account the Jacobian of the decoder in a VAE architecture.
@@ Appendix B Experimental Details
Environments.  We evaluate our methods on two control environments from the Gymnasiums library: Mountain Car Continuous and Reacher. The first is a classic control environment, which consists of a car placed stochastically in the middle of a sinusoidal valley, with the goal state on top of the right hill. The state is defined by two continuous variables: the position of the car along the $x$-axis $p\in[-1.2,0.6]$, and the velocity of the car $v\in[-0.07,0.07]$. The only possible action is to apply an acceleration $a\in[-1,1]$ to the car. The standard task is defined as $R_{\text{standard},t}=-0.1a^{2}$, until the goal is reached and a reward of $R_{\text{standard},t}=100$ is obtained, and the episode is ended. If the goal is not reached, the episode ends after 999 steps. We introduce three additional tasks: left, which is the same as the standard task, but with the goal moved to the top of the left hill ($p\leq-1.1$); height, which gives a reward of $R_{\text{height},t}=h^{2}$ at each time step that $h\geq 0.2$, with $h=\sin(3p)*0.45+0.55$ being the height of the car; speed, which gives a reward of $R_{\text{speed},t}=v^{2}$ at each time step. In the left task, the episode ends when the car reaches the left goal, while in the height and speed tasks, the episode ends when the car reaches the right goal.
@@
The second environment, Reacher, is a classic continuous control task consisting of a two-jointed robot arm, moving in a 2D space, with an end-effector called fingertip. The state is originally 10-dimensional, but we remove the coordinates of the target and the vector between the fingertip and the target. We end up with a 6-dimensional state composed of: $\cos(q_{1}),\cos(q_{2}),\sin(q_{1}),$ and $\sin(q_{2})$, the cosines and sines of the two join angles, and $\omega_{1}$ and $\omega_{2}$, their angular velocities. For the purpose of normalization, we consider the state bounded between the vectors [1,1,1,1,5,5][-1,-1,-1,-1,-5,-5] and $[1,1,1,1,5,5]$. The agent controls the arm by applying a distinct torque to each hinge, making the action space ${\mathcal{A}}=[-1,1]^{2}$. We disregard the standard task and instead define four new behavioral tasks that have the same reward shape $R_{\text{task}}=1$ if the condition is met, or 0 otherwise. In the speed task, the condition is that the linear velocity of the tip is greater than 6. In the clockwise and c-clockwise tasks, the condition is that the tangential velocity of the fingertip is greater than -11, or 1, respectively. Finally, in the radial task, the condition is that the radial velocity of the tip is greater than 3. The episodes terminate after 50 steps.
@@
Policies.  To model the policies, we use fully-connected, feed-forward, deterministic MLPs. Our choice of focusing on deterministic policies is dictated by the use of PGPE as optimization algorithm in the last stage, however, our pipeline is designed to be general. As such we believe there is apparent limitation to the use of stochastic policies instead. The input layer has $\mathcal{S}$} neurons, and it’s preceded by a normalization layer that standardizes the state features to have zero mean and unit variance. The hidden linear layers are followed by elu nonlinearities. The last layer has a $|{\mathcal{A}}|$ neurons, followed by a tanh activation to squash the action in the valid range. We test three different shapes of policies in Mountain Car: small policies composed of a single 4-neuron hidden layer; medium policies composed of two 32-neuron hidden layers; large policies composed of a 400-neuron hidden layer followed by a 300-neuron hidden layer. The number of parameters of the policies increases roughly by two orders of magnitude at each interval ($P_{\text{small}}=17$, $P_{\text{medium}}=1,185$, $P_{\text{large}}=121,801$). While in Reacher we use Medium policies composed of two 64-neuron hidden layers, with $P_{\text{Reacher}}=4738$.
@@
Policy Divergence.  To compute the divergence between policies, we instead estimate the distance of the deterministic actions of a subset of states. We consider the state spaces bounded as previously described, and we extract roughly $M=3000$ states. In Mountain Car we find them by discretizing the two dimensions and creating a grid, while in Reacher, we just sample them uniformly from the bounded state space. In the $k$-NN phase, we use $k=15$, and compute the distance between two policies as:
$D(\pi_{\bm{\theta}}||\pi_{{\bm{\theta}}^{\prime}})=\sqrt{\sum_{i=1}^{M}(\pi_{\bm{\theta}}(s_{i})-\pi_{{\bm{\theta}}^{\prime}}(s_{i}))^{2}}.$
While in the manifold learning phase we compute it as:
$D(\pi_{\bm{\theta}}||\pi_{{\bm{\theta}}^{\prime}})=\frac{1}{M}{\sum_{i=1}^{M}(\pi_{\bm{\theta}}(s_{i})-\pi_{{\bm{\theta}}^{\prime}}(s_{i}))^{2}}.$
@@
Autoencoder.  We use a simple, fully-connected, feed-forward, deterministic MLP to model the autoencoder. The shape of the autoencoder is the same for all the experiments: the input and output layers have size $P$, with the input layer being preceded by a standardization layer, and the output layer not being activated; the encoder has a 25-neuron hidden layer with an followed by a 10-neuron hidden layer; the decoder has the inverse shape of the encoder. The autoencoder is trained for 50 epochs using the Adam optimizer with an initial learning rate of 0.0001 and a batch size of 64. We employ a learning rate scheduler that halves the learning rate after 15 epochs of non-improvement, evaluated on a 20% random hold-out set. The empirical loss we train the autoencoder on is defined as:
${\mathcal{L}}_{B}=\frac{1}{N}\frac{1}{M^{\prime}}\sum_{i=1}^{N}\sum_{j=1}^{M^{\prime}}(\pi_{{\bm{\theta}}_{i}}(s_{j})-\pi_{\hat{\bm{\theta}}_{i}}(s_{j}))^{2},$
@@
where $N$ is the number of policies in the training dataset, and $M^{\prime}=1000$ is the dimension of the subset of the state set that we sample at each gradient step. In Mountain Car we set the latent dimension of the autoencoders to $k=1,2,3$, while in Reacher we use $k=1,2,3,5,8$.
@@
When we evaluate a latent space, we first compute the interquartile range for each dimension based on the spread of the training codes. Then, we discretize each dimension by an amount of points depending of the dimensions of the latent code: 100 points for 1D, 50 points for 2D, 17 points for 3D, 5 points for 5D, and 3 points for 8D. The decision is based solely on computational feasibility, and serves the purpose to have a rough conservative estimate of the range of encoded behaviors.
@@
Performance Recovery.  When comparing the policies found in the latent space with the ones belonging to the original dataset, we compute a behavior recovery metric in the following way. First, we average the dataset lower and upper bounds for all tasks the across three seeds with the same configuration, $lb_{D},ub_{D}$. Then we do the same among the discretized set of policies reconstructed from the latent space, $lb_{L},ub_{L}$. Finally, for each task, we compute the performance recovery as $\frac{ub_{L}-lb_{D}}{ub_{D}-lb_{D}}$. In Table 2, we provide the analysis for the reward left, which was omitted in Table 1.
@@
(Latent) PGPE.  As a byproduct of the low-dimensionality of the latent space, this framework is well suited to parameter-exploring PG methods. Algorithms like PGPE struggles with a high-dimensional set of parameters, such as those of a standard DRL network with hundreds of thousands of parameters. Yet, they can instead operate on the low-dimensional set of latent parameters while maintaining the expressivity of the original parameter space. As a bonus, the extension of PGPE to the latent space does not require computing the Jacobian of the decoder as in Eq. 8, as it can be seen as a deterministic addition to the black-box process that evaluates the parameters produced by the hyperpolicy $\nu_{\bm{\phi}}$. In fact, the objective defined in Eq. 2 can be rewritten under the latent PG formulation as $J^{R}({\bm{z}})=\mathbb{E}_{\tau\sim p(\cdot\mid{\bm{z}}),{\bm{z}}\sim\nu_{\bm{\phi}}}[R(\tau)]$, with the only change being that the probability density of the trajectories is given by the policy induced by the latent parameters as $p(\tau\mid{\bm{z}})=\mu(s_{0})\prod_{t=0}^{T}{\mathbb{P}}(s_{t+1}\mid s_{t},\pi_{g_{\bm{\zeta}}({\bm{z}})}(a_{t}|s_{t}))$. Finally, the gradient estimator at Eq. 3 is left unchanged, but for the change in parameter space from ${\bm{\theta}}_{i}$ to ${\bm{z}}_{i}$.
@@
We base our implementation of PGPE on an ask-and-tell implementation with symmetric sampling Toklu et al. (2020). We modify it to allow for numpy parallelization, reward normalization, center learning rate scheduling, learning $\log\sigma$ instead of $\sigma$ and natural gradient computation. The center is optimized with Adam, with momentum $0.2$. The log-standard deviation instead is learned through simple gradient ascent with fixed learning rate. For Mountain Car, we perform 75 seeded runs on 75 different autoencoders with the same hyperparameters: center learning rate 0.05, population size 4, initial standard deviation 0.6, standard deviation learning rate 0.1, and 50 generations. In Reacher, the learning rate has a linear annealing down to 20% of the initial value, and we use the same hyperparameters for all runs which are the following: center learning rate 0.01, population size 10, initial standard deviation 0.3, standard deviation learning rate 0.1, and 200 generations.

### v2

@@ 5.2 Supervised Fine-Tuning
(Q3) How can we fine-tune against specific tasks, leveraging the low-dimensional space? Does this offer any advantages?
To achieve this, we compared the effect of supervised fine-tuning on the latent space using a simple variant of PGPE, which we refer to as Latent PGPE, with various baselines, including PPO (Schulman et al., 2017), DDPG (Lillicrap et al., 2016), TD3 (Fujimoto et al., 2018), and SAC (Haarnoja et al., 2018). We also tested PGPE (Sehnke et al., 2008) in the high-dimensional parameter space, referred to as Parameter PGPE, as a sanity check. Our implementation of both Parameter PGPE and Latent PGPE can be found in Appendix B. All the algorithms were run in the best-performing configuration for the policy sizes. Interestingly, DRL baselines always struggle to optimize small policies and typically perform better with larger ones. On the contrary, Parameter PGPE benefits from having a reduced set of parameters to control; however, it generally suffers from high sample complexity. This evidence is reported in Appendix B (Fig. 7,8,9).
From these comparisons (reported in Fig. 4 and Fig. 5 for MC and RC, respectively), we were able to extract three main findings. First of all, the convergence rate and performance of Latent PGPE are positively correlated with the number of dimensions of the latent space, which confirms that larger latent spaces are indeed better shaped and with an easier optimization landscape, as hinted in Subsec. 5.1 as well. Secondly, Latent PGPE converges faster than all baselines in 7 out of 8 tasks, even though it does not always converge to the optimum. Finally, Latent PGPE is able to achieve comparable, if not better, performances to most of the baselines, even for complex tasks like speed in Fig. 4(c), 5(a). However, we observe that it fails to solve the height task in Fig. 4(d), due to the scarce representation of the high-performance policies in the unsupervised policy dataset.
Takeaways.  These results provide a positive answer to (Q3): Leveraging the learned low-dimensional representation of the behavioral manifold, the agent does not only achieve faster convergence, but better performance than state-of-the-art DRL algorithms in challenging sparse tasks. Unfortunately, the trade-off is that the fine-tuning performance is tied to the quality of the learned representation.
@@
Figure 5: Performance comparison in RC for different tasks. We report the average and 95% confidence interval over 10 runs. For clarity, the worst-performing baselines are omitted. A full study is reported in Appendix B.
@@ 6 Conclusions
We proposed a novel, unsupervised framework to address the sample inefficiency of Deep Reinforcement Learning by shifting the focus from parameter space to behavior space. Our approach successfully learns a compact latent manifold of policies, organized by behavioral similarity, using a generative model with an unsupervised behavioral reconstruction loss. Empirically, we showed that this approach can compress policy parameterizations by several orders of magnitude while preserving their functional expressivity. This compressed representation also allows for more efficient fine-tuning for downstream tasks via gradient-based optimization in the low-dimensional latent space.
@@
Ethics statement
This work presents fundamental research in reinforcement learning theory and algorithms. We have carefully reviewed the Code of Ethics and confirm that our research raises no ethical concerns.
@@
All experiments were run on 50 cores of an Intel(R) Xeon(R) Gold 64118H CPU, 1 TB of RAM, and one NVIDIA H100 GPU. The total wall-clock time to re-run all experiments is approximately 50 hours. Our proposed approach is detailed in Section 4 with full implementation details and hyperparameters included in Appendix B. The source code is available in the supplementary material and at github.com/DavideTenediniPoliMi/from-parameters-to-behaviors-unsupervised-compression-of-the-policy-space.
@@ Acknowledgments
We thank Andrea Fraschini for his assistance in conducting the baseline experiments. This work was supported by the Italian Ministry of University and Research (MUR) under the National Recovery and Resilience Plan (NRRP), and by the European Union (EU) under the NextGenerationEU project (CUP D43C24001670008).
Appendix A Related Works
@@
Simplification of the Policy Space.  A variety of works (among others, Gregor et al., 2017; Eysenbach et al., 2018; Achiam et al., 2018; Hansen et al., 2019) have proposed methods to simplify the policy space. Yet, those policies should be generally intended as mere initializations for supervised fine-tuning, which falls back to operating in the original policy space once the downstream task is revealed. To the best of our knowledge, the only other work defining a formal criterion to operate a compression of the policy space is Mutti et al. (2022). Yet, this paper seeks a way to reduce the cardinality of the policy space, rather than its dimensionality. Moreover, the constraints defining a valid compression are stricter than ours, resulting in an optimization problem that is known to be NP-hard. Finally, their work does not provide a way to perform supervised fine-tuning in a scalable way.
Weight Space Learning  The goal of our work, namely, to learn a latent representation of neural network parameters, is shared by the field of Weight Space Learning (WSL). In RL, recent works have investigated task-specific generative models, conditioning them on goals (Faccio et al., 2023) or performance checkpoints (Peebles et al., 2022), unlike our fully unsupervised approach. Our approach is more in line with the works on hyper-representations (Schürholt et al., 2021; 2022; 2024) of networks trained for supervised tasks. Hyper-representations are task-agnostic low-dimensional embeddings of neural networks learned from a zoo of trained models. Beyond its novel application to RL, our approach differs from these works in two ways. First, our “zoo” does not require trained experts. Second, and more critically, WSL methods must often contend with the vast number of parameter-space symmetries (e.g., neuron permutations and scaling (Kůrková & Kainen, 1994)) using complex architectures or workarounds. Our behavioral reconstruction loss fundamentally sidesteps this challenge. By compressing policies based on their function rather than their specific parameters, our autoencoder naturally assimilates these redundant parameterizations. This concept of functional compression is similar to “policy fingerprinting” (Faccio et al., 2023), which also gauges behavior independent of weights.
@@
Policy Manifold and Quality Diversity.  The idea of employing generative models to learn a compressed representation of the policy space has received some recent attention outside of the Weight Space Learning field. Rakicevic et al. (2021) hypothesized that there might be a low-dimensional manifold embedded in the policy parameter space, even if they did not characterize it formally. Chang et al. (2019) trains a Variational AE to reconstruct the weights of pre-trained expert policies to learn expert-agent embeddings and analyze the latent structure of the solution space. A similar architecture has also been applied in the field of Quality Diversity, either to improve the sample efficiency of diversity-based search Rakicevic et al. (2021) or to distill a large policy archive into a compact generative model  (Hegde et al., 2023). Notably, all the methods above employ VAE architectures with a parameter-reconstruction loss, which allows only moderate compression ratios of up to $19:1$ (Hegde et al., 2023). In comparison, this paper introduces a fully unsupervised pipeline that focuses on compressing a behavioral loss, intending to provide a compact space for latent policy optimization as well while retaining greater compression abilities.
@@
Policy Optimization.  First-order methods have been extensively employed to address PO (Peters & Schaal, 2008; Lillicrap et al., 2016) as well as natural gradients (Kakade, 2001) and trust-region methods (Schulman et al., 2017). Yet in this work, we built on the long tradition of PGPE Algorithms (Sehnke et al., 2008; Rückstiess et al., 2010; Miyamae et al., 2010; Montenegro et al., 2024), as their hard scalability to large parameter spaces is notoriously a blocking factor. Finally, we notice that Rakicevic et al. (2021) indeed proposed a method to optimize the diversity of the policies by taking into account the Jacobian of the decoder in a VAE architecture.
@@ Appendix B Experimental Details
Environments.  We evaluate our methods on four control environments from the Gymnasium library: Mountain Car Continuous, Reacher, Hopper, and HalfCheetah. The first is a classic control environment, which consists of a car placed stochastically in the middle of a sinusoidal valley, with the goal state on top of the right hill. The state is defined by two continuous variables: the position of the car along the $x$-axis $p\in[-1.2,0.6]$, and the velocity of the car $v\in[-0.07,0.07]$. The only possible action is to apply an acceleration $a\in[-1,1]$ to the car. The standard task is defined as $R_{\text{standard},t}=-0.1a^{2}$, until the goal is reached and a reward of $R_{\text{standard},t}=100$ is obtained, and the episode ends. If the goal is not reached, the episode ends after 999 steps. We introduce three additional tasks: left, which is the same as the standard task, but with the goal moved to the top of the left hill ($p\leq-1.1$); height, which gives a reward of $R_{\text{height},t}=h^{2}$ at each time step for which $h\geq 0.2$, with $h=\sin(3p)*0.45+0.55$ being the height of the car; speed, which gives a reward of $R_{\text{speed},t}=v^{2}$ at each time step. In the left task, the episode ends when the car reaches the left goal, while in the height and speed tasks, the episode ends when the car reaches the right goal.
@@
The second environment, Reacher, is a classic continuous control task consisting of a two-jointed robot arm, moving in a 2D space, with an end-effector called fingertip. The state is originally 10-dimensional, but we remove the coordinates of the target and the vector between the fingertip and the target. We end up with a 6-dimensional state composed of: $\cos(q_{1}),\cos(q_{2}),\sin(q_{1}),$ and $\sin(q_{2})$, the cosines and sines of the two joint angles, and $\omega_{1}$ and $\omega_{2}$, their angular velocities. For the purpose of normalization, we consider the state bounded between the vectors [1,1,1,1,5,5][-1,-1,-1,-1,-5,-5] and $[1,1,1,1,5,5]$. The agent controls the arm by applying a distinct torque to each hinge, making the action space ${\mathcal{A}}=[-1,1]^{2}$. We disregard the standard task and instead define four new behavioral tasks that have the same reward shape $R_{\text{task}}=1$ if the condition is met, or 0 otherwise. In the speed task, the condition is that the linear velocity of the tip is greater than 6. In the clockwise and c-clockwise tasks, the condition is that the tangential velocity of the fingertip is greater than -11, or 1, respectively. Finally, in the radial task, the condition is that the radial velocity of the tip is greater than 3. The episodes terminate after 50 steps.
The third environment, Hopper, is a classic continuous control task that models a 1-legged robot with 3 joints. The state space is 11-dimensional, capturing the robot’s $z$-position, joint angles, and corresponding velocities. For the purpose of normalization, we consider the state bounded between the vectors [0.7,0.2,2.7,2.7,0.8,5.0,5.0,5.0,5.0,5.0,5.0][0.7,-0.2,-2.7,-2.7,-0.8,-5.0,-5.0,-5.0,-5.0,-5.0,-5.0] and [1.5,0.2,0.0,0.0,0.8,5.0,5.0,5.0,5.0,5.0,5.0][1.5,0.2,0.0,0.0,0.8,5.0,5.0,5.0,5.0,5.0,5.0]. The agent controls the arm by applying a distinct torque to each hinge, making the action space ${\mathcal{A}}=[-1,1]^{3}$. Analogously to Reacher, we disregard the standard task and instead define four new behavioral tasks that have the same reward shape $R_{\text{task}}=1$ if the condition is met, or 0 otherwise. In the forward, backward, and standstill tasks, the condition is that the $x$-axis velocity is greater than 1, lower than -1, or between -0.05 and 0.05. In the jump task, the $z$ position has to be greater than 1.3. The episodes terminate after 1000 steps.
The fourth environment, HalfCheetah, is a classic continuous control task that models a 2D, two-legged robot with 6 actuated joints. The state space is 18-dimensional, including the robot’s body position, rotational angles, and joint velocities. For the purpose of normalization, we consider the state bounded between the vectors $0,\allowbreak-\pi,\allowbreak-0.52,-0.785,\allowbreak-0.4,\allowbreak-1,\allowbreak-1.2,\allowbreak-0.5,\allowbreak-5.0,-5.0,-5.0,\allowbreak-5.0,-5.0,-5.0,\allowbreak-5.0,-5.0,\allowbreak-5.0]$ and $1.5,\pi,1.05,\allowbreak 0.785,0.785,\allowbreak 0.7,0.87,0.5,\allowbreak 5.0,5.0,5.0,\allowbreak 5.0,5.0,\allowbreak 5.0,5.0,5.0,\allowbreak 5.0]$. The agent controls the arm by applying a distinct torque to each hinge, making the action space ${\mathcal{A}}=[-1,1]^{6}$. The tasks are almost the same as for Hopper. In the forward and backward, the condition is that the $x$-axis velocity is greater than 2, or lower than -2. In the frontflip and backflip, the task condition is that the cumulative rotation along the $y$-axis increases or decreases by a full rotation from the initial angle. The episodes terminate after 1000 steps.
@@
Policies.  To model the policies, we use fully-connected, feed-forward, deterministic MLPs. Our choice of focusing on deterministic policies is dictated by the use of PGPE as an optimization algorithm in the last stage; however, our pipeline is designed to be general. As such, we believe there is no apparent limitation to the use of stochastic policies instead. The input layer has $\mathcal{S}$} neurons and is preceded by a normalization layer that standardizes the state features to have zero mean and unit variance. The hidden linear layers are followed by elu nonlinearities. The last layer has $|{\mathcal{A}}|$ neurons, followed by a tanh activation to squash the action into the valid range. We test three different shapes of policies in Mountain Car: Small policies composed of a single 4-neuron hidden layer; Medium policies composed of two 32-neuron hidden layers; Large policies composed of a 400-neuron hidden layer followed by a 300-neuron hidden layer. The number of parameters of the policies increases roughly by two orders of magnitude at each interval ($P_{\text{Small}}=17$, $P_{\text{Medium}}=1,185$, $P_{\text{Large}}=121,801$). For Hopper and HalfCheetah, we use the same shape of Medium policies. While in Reacher we use Medium policies composed of two 64-neuron hidden layers, with $P_{\text{Reacher}}=4738$.
@@
Policy Divergence.  To compute the divergence between policies, we instead estimate the distance of the deterministic actions over a subset of states. We consider the state spaces bounded as previously described, and we extract roughly $M=3000$ states. In Mountain Car, we find them by discretizing the two dimensions and creating a grid, while in MuJoCo environments, we simply sample them uniformly from the bounded state space. In the $k_{n}$-NN phase, we use $k_{n}=15$, and compute the distance between two policies as:
$D(\pi_{\bm{\theta}}\parallel\pi_{{\bm{\theta}}^{\prime}})=\sqrt{\sum_{i=1}^{M}\left\|\pi_{\bm{\theta}}(s_{i})-\pi_{{\bm{\theta}}^{\prime}}(s_{i})\right\|^{2}_{2}}.$
While in the manifold learning phase, we compute it as:
$D(\pi_{\bm{\theta}}\parallel\pi_{{\bm{\theta}}^{\prime}})=\frac{1}{M}{\sum_{i=1}^{M}\left\|\pi_{\bm{\theta}}(s_{i})-\pi_{{\bm{\theta}}^{\prime}}(s_{i})\right\|^{2}_{2}}.$
@@
Autoencoder.  We use a simple, fully-connected, feed-forward, deterministic MLP to model the autoencoder. The shape of the autoencoder is the same for all the experiments. The input and output layers have size $P$, with the input layer being preceded by a standardization layer, and the output layer not being activated; the encoder has a 25-neuron hidden layer followed by a 10-neuron hidden layer; the decoder has the mirrored shape of the encoder. The first layer of the encoder and the first layer of the decoder are followed by elu nonlinearities. The autoencoder is trained for 50 epochs using the Adam optimizer with an initial learning rate of 0.0001 and a batch size of 64. We employ a learning rate scheduler that halves the learning rate after 15 epochs of non-improvement, evaluated on a 20% random hold-out set. The empirical loss used to train the autoencoder is defined as:
${\mathcal{L}}_{B}=\frac{1}{N}\frac{1}{M^{\prime}}\sum_{i=1}^{N}\sum_{j=1}^{M^{\prime}}\left\|\pi_{{\bm{\theta}}_{i}}(s_{j})-\pi_{\hat{\bm{\theta}}_{i}}(s_{j})\right\|^{2}_{2},$
@@
where $N$ is the number of policies in the training dataset, $M^{\prime}=1000$ is the size of the subset of the state set that we sample at each gradient step, and $\pi_{\hat{\bm{\theta}}_{i}}$ is the reconstructed policy. In Mountain Car, we set the latent dimension of the autoencoders to $k=1,2,3$, while in Reacher, we use $k=1,2,3,5,8$. In Hopper and HalfCheetah, $k=5,8,16$.
@@
When we evaluate a latent space, we first compute the interquartile range for each dimension based on the spread of the training codes. Then, we discretize each dimension by a variable number of points depending on the dimension of the latent code: 100 points for 1D, 50 points for 2D, and 17 points for 3D. For more than three latent dimensions, we sample 10000 policies from the same bounded latent space at random. The decision is based solely on computational feasibility and serves the purpose of having a rough conservative estimate of the range of encoded behaviors.
@@
Performance Recovery.  When comparing the policies found in the latent space with the ones belonging to the original dataset, we compute a behavior recovery metric in the following way. First, we average the dataset lower and upper bounds for all tasks across three seeds with the same configuration, $lb_{D},ub_{D}$. Then we do the same for the discretized set of policies reconstructed from the latent space, $lb_{L},ub_{L}$. Finally, for each task, we compute the performance recovery as $\frac{ub_{L}-lb_{D}}{ub_{D}-lb_{D}}$. In Table 3, we provide the analysis for the reward left, which was omitted in Table 1.
@@
(Latent) PGPE.  As a byproduct of the low-dimensionality of the latent space, this framework is well suited to parameter-exploring PG methods. Algorithms like PGPE struggle with a high-dimensional set of parameters, such as those of a standard DRL network with hundreds of thousands of parameters. Yet, they can instead operate on the low-dimensional set of latent parameters while maintaining the expressivity of the original parameter space. As a bonus, the extension of PGPE to the latent space does not require computing the Jacobian of the decoder as in Eq. 5, as it can be seen as a deterministic addition to the black-box process that evaluates the parameters produced by the Gaussian hyperpolicy $\nu_{\bm{\phi}}$, where ${\bm{\phi}}=({\bm{\mu}},{\bm{\sigma}})$ is the vector of means and standard deviations parameterizing the Gaussian distributions over the latent parameters. In fact, the objective defined in Eq. 2 can be rewritten under the latent PG formulation as $J^{R}({\bm{z}})=\mathbb{E}_{\tau\sim p(\cdot\mid{\bm{z}}),{\bm{z}}\sim\nu_{\bm{\phi}}}[R(\tau)]$, with the only change being that the probability density of the trajectories is given by the policy induced by the latent parameters as $p(\tau\mid{\bm{z}})=\mu(s_{0})\prod_{t=0}^{T}{\mathbb{P}}(s_{t+1}\mid s_{t},a_{t})\ \pi_{g_{\bm{\zeta}}({\bm{z}})}(a_{t}\mid s_{t})$. Finally, the gradient estimator at Eq. 3 is left unchanged, but for the change in parameter space from ${\bm{\theta}}_{i}$ to ${\bm{z}}_{i}$.
@@
We base our implementation of PGPE on an ask-and-tell implementation with symmetric sampling (Toklu et al., 2020). We modify it to allow for numpy parallelization, reward normalization, center learning rate scheduling, learning $\log\sigma$ instead of $\sigma$, and natural gradient computation. The center is optimized with Adam, with momentum $0.2$. The log-standard deviation instead is learned through simple gradient ascent with fixed learning rate. For Mountain Car, we perform 75 seeded runs on 75 different autoencoders with the same hyperparameters: center learning rate 0.05, population size 4, initial standard deviation 0.6, standard deviation learning rate 0.1, and 50 generations. In Reacher, the learning rate has a linear annealing down to 20% of the initial value, and we use the same hyperparameters for all runs, which are the following: center learning rate 0.1, population size 10, initial standard deviation 0.3, standard deviation learning rate 0.1, and 200 generations. In both cases, each sample of the population is evaluated on a single episode.
Given the episodic nature of PGPE, each generation can take an arbitrary number of environment steps to evaluate the samples of the population. Since the center is changed only once at the end of a generation, we evaluate the learning curve at different sample checkpoints, contrary to the standard StableBaselines3 approach. This causes different runs to have different evaluation checkpoints ($x$-axis of the learning curve) and a different number of total samples used. To visualize an aggregate learning curve across multiple runs, we take two measures: we interpolate the various curves along the $x$-axis; we extend the final evaluation of each run along the $x$-axis so that all runs have identical lengths, with the assumption that the algorithm has already converged by the end of the training. To ensure the stability of the final evaluation, we evaluate the center solution on 100 episodes (instead of 10 episodes used during training).